# Quantum Thermal Analogs of Electric Circuits: A Universal Approach

### v3

@@ Abstract
We model a quantum thermal resistor identified as the strength of interaction Hamiltonian to find the heat current between two or more qubits. We establish the laws of a quantum thermal analog of an electric circuit, analogous to Kirchhoff’s current and voltage laws, treating a qubit as the node. We also prove the existence of a quantum thermal step transformer. We develop two novel quantum thermal circuits—quantum thermal super Wheatstone bridge, and a thermal adder circuit (motivated by the operational amplifier circuit, where the voltage output is the algebraic sum of all the inputs)—and use the above framework to study them. It is shown that the thermal diode, transistor, and the standard Wheatstone bridge circuits fall into the framework constructed here, the details of which are presented in Supplemental Material [54]. The framework introduced is the first of its kind, providing a fundamental tool to build a comprehensive quantum thermal circuit theory. The fact that many thermal versions of important classical electric circuits can be studied in a unified manner under the umbrella of this framework will interest a broader audience in the juxtaposition of quantum thermodynamics and electric circuits.
@@
Quantum thermal voltage—We now develop the quantum analog of the thermal voltage driving the thermal current. To this end, we determine the effective temperature of qubits 1 and 2 in the steady state. The reduced state of qubit 1 from the steady state, Eq. 2, is given by $\rho^{SS}_{1}=\left(\rho^{SS}_{00}+\rho^{SS}_{11}\right)\ket{0}\bra{0}+\left(\rho^{SS}_{22}+\rho^{SS}_{33}\right)\ket{1}\bra{1}.$ | . The effective temperature $T_{1}$ of the qubit 1 is found by comparing the state $\rho^{SS}_{1}$ with the state $e^{-H_{1}/T_{1}}/{\rm Tr}\left(e^{-H_{1}/T_{1}}\right)$, where $H_{1}=\frac{\omega_{1}}{2}\sigma^{z}_{1}$, leading to $T_{1}=\omega_{1}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{11}}-1\right)=\omega_{1}/\log\left(\frac{\gamma_{B}\eta_{00}}{x_{0}\alpha_{00}}-1\right).$ Note that $T_{1}$ can be envisaged as the thermal potential difference between qubit 1 and bath $A$, that is, $V_{1A}=T_{1}-0=T_{1}$. This potential difference drives the thermal current $\mathcal{J}_{A1}$, which explicitly in terms of $T_{1}$ can be written as $\mathcal{J}_{A1}=\frac{-\gamma_{A}\omega_{1}}{1+e^{\omega_{1}/T_{1}}}.$ In the low and high-temperature limits, this is
@@
Figure 2: Variation of the thermal current $|\mathcal{J}_{A1}|$ between the qubit 1 and the bath $A$ as a function of the thermal potential difference $V_{1A}=T_{1}$ for $\gamma_{A}=0.5$ and $\omega_{1}=1.0$.
@@
We use the above framework to develop novel circuits, to which we coin the names quantum thermal super Wheatstone bridge and quantum thermal adder. Further, thermal devices discussed in recent times, for example, a Wheatstone bridge, diode, and quantum thermal transistor, are presented using our framework in [54].
@@
In condition 2, we explore the scenario where the center qubit 7 is not affected by the thermal transport around it. To do so, we allow the input and output of heat from single connections to the square made by qubits 3, 4, 5, and 6, thereby making $J_{23}=J_{61}=0$. Further, we allow only two connections to qubit $7$, for example, with qubit 3 and 6, implying $J_{57}=J_{47}=0$. The balance condition is obtained when $J_{34}=J_{46}$ and $J_{35}=J_{56}$ ($J_{34}$ need not be equal to $J_{35}$), together with $\omega_{3}=\omega_{6}=\omega_{7}$. In this condition, see Fig. 6(b), the currents $\mathcal{J}_{37}=\mathcal{J}_{67}=0$, making the qubit 7 unaffected by the thermal transport around it. Interestingly, if we increase the number of qubits between qubits 3 and 6, keeping their $\omega_{i}$’s equal, they will also remain unaffected by the thermal transport around them [54]. Condition 3 of the super Wheatstone bridge is identical to the Wheatstone bridge condition [54]. We can form many Wheatstone bridges inside this super Wheatstone bridge; for example, see the black dashed line between qubits 1, 5, 7, and 6. Here, we take $J_{35}=J_{37}=J_{46}=0$ to form a particular Wheatstone bridge. The balance condition is achieved when $J_{15}=J_{16}$ and $J_{57}=J_{67}$ ($J_{15}$ need not be equal to $J_{57}$); and then the current between qubits 5 and 6, $\mathcal{J}_{56}$, becomes zero, see Fig. 6(c), a classic Wheatstone bridge condition.
@@
This thermal circuit is motivated by the DC operational amplifier’s adder circuit, where the output voltage of the device is the sum of all the input voltages. Here, the qubits $1,2,...,n$ interact with the baths $I,II,...,N$, at temperatures $T_{I},T_{II},...,T_{N}$, respectively. The effective temperatures of the qubits $1,2,...,n$ are $T_{1},T_{2},...,T_{n}$, respectively. The thermal potential difference between qubit 1 and bath $I$ is $V_{I1}=T_{I}-T_{1}$ and similarly for the other qubits. Qubits $1,2,...,n$ also interact with qubit $\alpha$ connected to a zero temperature bath $T_{A}$ and the thermal potential difference between qubit $\alpha$ and bath $A$ is $V_{\alpha A}=T_{\alpha}-T_{A}=T_{\alpha}$, where $T_{\alpha}$ is the effective temperature of qubit $\alpha$. The Hamiltonian of the system is $H_{add}=\frac{\omega_{\alpha}}{2}\sigma^{z}_{\alpha}+\sum_{i=1}^{n}\left[\frac{\omega_{i}}{2}\sigma^{z}_{i}+J_{i\alpha}\left(\sigma^{x}_{\alpha}\sigma^{x}_{i}+\sigma^{y}_{\alpha}\sigma^{y}_{i}\right)\right]$. Using the thermal current law, $\mathcal{J}_{A\alpha}=\sum_{k=1}^{n}\mathcal{J}_{k\alpha}$ and Eq. (5), we can write $\mathcal{J}_{A\alpha}=-\sum_{k=1}^{n}\frac{\omega_{\alpha}}{\omega_{k}}\mathcal{J}_{\alpha k}$. But $\mathcal{J}_{\alpha k}=\mathcal{J}_{jk}$ (for $(j,k)=(I,1),(II,2),...(N,n)$), implying the following relation
@@
which for $\omega_{k}=\omega_{\alpha}\forall k$ results into $\mathcal{J}_{A\alpha}=-\left(\mathcal{J}_{I1}+\mathcal{J}_{II2}+...+\mathcal{J}_{Nn}\right)$. The above equation exactly mimics the relation between the input and output voltages of an adder circuit. The currents $\mathcal{J}_{jk}$ and $\mathcal{J}_{A\alpha}$ are related to the effective temperature of the corresponding qubits $T_{k}$ via $\mathcal{J}_{jk}=\gamma_{j}\omega_{k}\left[-1+\coth\left(\frac{\omega_{k}}{2T_{j}}\right)\tanh\left(\frac{\omega_{k}}{2T_{k}}\right)\right],$ for $j=A,I,II,...N$ and $k=\alpha,1,2,...n$. Taking $\gamma_{A}=\gamma_{I}=...=\gamma_{N}$, $\omega_{1}=\omega_{2}=...=\omega_{n}=\omega$, $J_{1\alpha}=J_{2\alpha}=...=J_{n\alpha}$ and $T_{I}=T_{II}=...=T_{N}=T$, we get a simplified relation between the effective temperatures $T_{x}$ for $x=1,2,...,n$ and the qubit $\alpha$ ($T_{\alpha}$)
@@
Conclusion—We have presented a comprehensive schematic for a quantum thermal circuit theory, paving a roadmap to construct quantum thermal integrated circuits, the important ingredients of which are quantum thermal diodes and transistors. We prove the existence of Kirchhoff-like voltage and current laws, with interaction strength playing the role of circuit resistance. We further demonstrate how heat currents can be manipulated to construct a quantum step transformer. Using this, we develop a quantum super Wheatstone bridge with many interesting balance conditions, also containing the standard quantum Wheatstone bridge, where unknown Hamiltonian strengths can be determined by the profiling of heat currents. This circuit can have a plethora of applications in process tomography, metrology, and quantum sensing. A model of a quantum thermal adder circuit for temperature gradients was also constructed, creating further opportunities to design thermal operational amplifiers and other integrated circuits. We also bring out the interrelation between the heat current developed here and the established forms of current used in the literature. Thus, for example, the spin current, which can be experimentally measured [60, 61, 53], provides a handle for experimental determination of the heat current. The framework of the circuits developed here can be adapted to experimental realization. Thus, for example, progress has been made towards the diode architecture with pentamethyl-disilane in an NMR register [62]. Also, an experimental study of thermodynamic uncertainty relation in the two-qubit NMR spin system, sodium fluorophosphate, was performed in [63], and an experimental setup of a quantum thermal transistor using a superconducting circuit was proposed in [64]. Along similar lines, experimental demonstration of other quantum thermal circuits discussed here could be envisaged. Our analysis opens up several research avenues, for example, studies in thermal conductivity, balance conditions of the super Wheatstone bridge, and other analogs of electric circuits. It is thus sufficient to say that the present work offers the fundamentals of a comprehensive theory of quantum thermal analogs of electric circuits, paving the way for the corresponding integrated circuits.
@@ II Details on the forms of heat currents and proof of quantum thermal version of Kirchhoff’s current law
Consider a general system of $n$ nodes (labeled $1$, $2$, $3$, …, $n$), where each node is coupled to all other nodes. Further, each node is weakly coupled to its respective bosonic thermal bath (labeled $I$, $II$, $III$, …, $N$). The system’s Hamiltonian in this setup is given by
@@
where $\mathcal{D}_{jk}(\rho)=\gamma_{j}(\tilde{N}_{jk}+1)\left(L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\left\{L_{k}^{\dagger}L_{k},\rho\right\}\right)+\gamma_{j}\tilde{N}_{jk}\left(L_{k}^{\dagger}\rho L_{k}-\frac{1}{2}\left\{L_{k}L_{k}^{\dagger},\rho\right\}\right)$, and $\tilde{N}_{jk}=\frac{1}{e^{\beta_{j}\omega_{k}}-1}$, with $\beta_{j}=T_{j}^{-1}$ for $(j,k)=(I,1),(II,2),...,(N,n)$. Further, $L_{k}$ are the Lindblad jump operators, which in the case of qubits are $\sigma^{-}_{k}$. Let us pick a node $m$ (Hamiltonian $H_{m}$) from the $n$ nodes. This node $m$ would have quantum heat currents from all the sources, that is, from all the other nodes that interact with it, as well as from the bosonic bath it is in interaction with (let the corresponding bath be $M$). The time derivative of the expectation value of $H_{m}$ is
@@
It can be easily verified from the above equation that ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]=0$ for $(j,k)=(I,1),(II,2),...,(N,n)$ except when $j=M$ and $k=m$. Therefore, Eq. (18) reduces to
@@
The commutator of $H_{m}$ with $H_{S}$ can further be simplified to give $[H_{m},H_{S}]=-\sum_{l=1,l\neq m}^{n}[H_{lm},H_{m}]$, using which we can write
@@ II.1 Comparison of the derived heat current with the established forms in the literature
where $j_{k-1,k}$ is the flux incoming from the $k-1$-th qubit to $k$-th qubit and similarly $j_{k,k+1}$ is the flux outgoing from the $k$-th qubit to $k+1$-th qubit. The left-hand side in the above equation can be obtained using the master equation (15) as
@@
where the terms ${\rm Tr}\left[\sigma^{z}_{k}\mathcal{D}_{A1}\right]$ and ${\rm Tr}\left[\sigma^{z}_{k}\mathcal{D}_{B2}\right]$ become zero for $k\neq 1,2$. Further, the commutator $[H_{S},\sigma^{z}_{k}]$ simplifies to $[H_{S},\sigma^{z}_{k}]=[H_{k-1,k},\sigma^{z}_{k}]+[H_{k,k+1},\sigma^{z}_{k}]$, where $H_{k-1,k}$ and $H_{k,k+1}$ are the interaction Hamiltonians between qubits at sites $k-1$ and $k$ and between qubits at sites $k$ and $k+1$, respectively. This follows from the fact that all other terms in $H_{S}$ commute with $\sigma^{z}_{k}$. Now, the above equation can be rewritten as
@@
$j_{k,k+1}=2J\left\langle\sigma^{x}_{k}\sigma^{x}_{k+1}-\sigma^{y}_{k}\sigma^{y}_{k+1}\right\rangle=I_{M}.$ (34)
A form of heat current is also found by multiplying the spin current with the frequency of the qubit [58]. However, on comparing the spin current $I_{M}$ and the form of heat current obtained in Eq. (23), we find the magnitudes of both are related by
@@
The forms of heat current, Eqs. (23) and (24), are more general in the sense that the expressions of heat current, along with their correct frequency scaling factors, can be obtained in scenarios where there are chains of harmonic oscillators or a mixed chain of harmonic oscillators and spins and for arbitrary interactions between them. Further, there is an advantage in using these forms; for example, in a complicated scenario, such as a circuit of a thermal super Wheatstone bridge, the magnitude and direction of the heat current can be exactly determined.

### v4

@@ Abstract
We model a quantum thermal resistor identified by the strength of the interaction Hamiltonian to find the heat current between two or more qubits. We establish the laws of a quantum thermal analog of an electric circuit, analogous to Kirchhoff’s current and voltage laws, treating a qubit as the node. We also prove the existence of a quantum thermal step transformer. We develop two novel quantum thermal circuits—a quantum thermal super Wheatstone bridge, and a thermal adder circuit (motivated by the operational amplifier circuit, where the voltage output is the algebraic sum of all the inputs)—and use the above framework to study them. It is shown that the thermal diode, transistor, and the standard Wheatstone bridge circuits fall into the framework constructed here, the details of which are presented in Supplemental Material [54]. The framework introduced is the first of its kind, providing a fundamental tool to build a comprehensive quantum thermal circuit theory. The fact that many thermal versions of important classical electric circuits can be studied in a unified manner under the umbrella of this framework will interest a broader audience in the juxtaposition of quantum thermodynamics and electric circuits.
@@
[…]
@@
Figure 2: Variation of the thermal current $|\mathcal{J}_{A1}|$ between the qubit 1 and the bath AA as a function of the thermal potential difference $V_{1A}=T_{1}$ for $\gamma_{A}=0.5$ and $\omega_{1}=1.0$.
@@
We use the above framework to develop novel circuits, to which we coin the names quantum thermal super Wheatstone bridge and quantum thermal adder. Further, thermal devices discussed in recent times, for example, a Wheatstone bridge, a diode, and a quantum thermal transistor, are presented using our framework in [54].
@@
In condition 2, we explore the scenario where the center qubit 7 is unaffected by the thermal transport around it. To do so, we allow the input and output of heat from single connections to the square made by qubits 3, 4, 5, and 6, thereby making $J_{23}=J_{61}=0$. Further, we allow only two connections to qubit $7$, for example, with qubit 3 and 6, implying $J_{57}=J_{47}=0$. The balance condition is obtained when $J_{34}=J_{46}$ and $J_{35}=J_{56}$ ($J_{34}$ need not be equal to $J_{35}$), together with $\omega_{3}=\omega_{6}=\omega_{7}$. In this condition, see Fig. 6(b), the currents $\mathcal{J}_{37}=\mathcal{J}_{67}=0$, making the qubit 7 unaffected by the thermal transport around it. Interestingly, if we increase the number of qubits between qubits 3 and 6, keeping their $\omega_{i}$’s equal, they will also remain unaffected by the thermal transport around them [54]. Condition 3 of the super Wheatstone bridge is identical to the Wheatstone bridge condition [54]. We can form many Wheatstone bridges inside this super Wheatstone bridge; for example, see the black dashed line between qubits 1, 5, 7, and 6. Here, we take $J_{35}=J_{37}=J_{46}=0$ to form a particular Wheatstone bridge. The balance condition is achieved when $J_{15}=J_{16}$ and $J_{57}=J_{67}$ ($J_{15}$ need not be equal to $J_{57}$); and then the current between qubits 5 and 6, $\mathcal{J}_{56}$, becomes zero, see Fig. 6(c), a classic Wheatstone bridge condition.
@@
This thermal circuit is motivated by the DC operational amplifier’s adder circuit, where the output voltage of the device is the sum of all the input voltages. Here, the qubits 1,2,,n1,2,...,n interact with the baths I,II,,NI,II,...,N, at temperatures $T_{I},T_{II},...,T_{N}$, respectively. The effective temperatures of the qubits 1,2,,n1,2,...,n are $T_{1},T_{2},...,T_{n}$, respectively. The thermal potential difference between qubit 1 and bath $I$ is $V_{I1}=T_{I}-T_{1}$ and similarly for the other qubits. Qubits 1,2,,n1,2,...,n also interact with qubit $\alpha$ connected to a zero temperature bath $T_{A}$ and the thermal potential difference between qubit $\alpha$ and bath $A$ is $V_{\alpha A}=T_{\alpha}-T_{A}=T_{\alpha}$, where $T_{\alpha}$ is the effective temperature of qubit $\alpha$. The Hamiltonian of the system is $H_{add}=\frac{\omega_{\alpha}}{2}\sigma^{z}_{\alpha}+\sum_{i=1}^{n}\left[\frac{\omega_{i}}{2}\sigma^{z}_{i}+J_{i\alpha}\left(\sigma^{x}_{\alpha}\sigma^{x}_{i}+\sigma^{y}_{\alpha}\sigma^{y}_{i}\right)\right]$. Using the thermal current law, $\mathcal{J}_{A\alpha}=\sum_{k=1}^{n}\mathcal{J}_{k\alpha}$ and Eq. (5), we can write $\mathcal{J}_{A\alpha}=-\sum_{k=1}^{n}\frac{\omega_{\alpha}}{\omega_{k}}\mathcal{J}_{\alpha k}$. But $\mathcal{J}_{\alpha k}=\mathcal{J}_{jk}$ (for (j,k)=(I,1),(II,2),(N,n)(j,k)=(I,1),(II,2),...(N,n)), implying the following relation
@@
which for $\omega_{k}=\omega_{\alpha}\forall k$ results into $\mathcal{J}_{A\alpha}=-\left(\mathcal{J}_{I1}+\mathcal{J}_{II2}+...+\mathcal{J}_{Nn}\right)$. The above equation exactly mimics the relation between the input and output voltages of an adder circuit. The currents $\mathcal{J}_{jk}$ and $\mathcal{J}_{A\alpha}$ are related to the effective temperature of the corresponding qubits $T_{k}$ via $\mathcal{J}_{jk}=\gamma_{j}\omega_{k}\left[-1+\coth\left(\frac{\omega_{k}}{2T_{j}}\right)\tanh\left(\frac{\omega_{k}}{2T_{k}}\right)\right],$ for j=A,I,II,Nj=A,I,II,...N and $k=\alpha,1,2,...n$. Taking $\gamma_{A}=\gamma_{I}=...=\gamma_{N}$, $\omega_{1}=\omega_{2}=...=\omega_{n}=\omega$, $J_{1\alpha}=J_{2\alpha}=...=J_{n\alpha}$ and $T_{I}=T_{II}=...=T_{N}=T$, we get a simplified relation between the effective temperatures $T_{x}$ for x=1,2,,nx=1,2,...,n and the qubit $\alpha$ ($T_{\alpha}$)
@@
Conclusion—We have presented a comprehensive schematic for a quantum thermal circuit theory, paving a roadmap to construct quantum thermal integrated circuits, the important ingredients of which are quantum thermal diodes and transistors. We prove the existence of Kirchhoff-like voltage and current laws, with interaction strength playing the role of circuit resistance. We further demonstrate how heat currents can be manipulated to construct a quantum step transformer. Using this, we develop a quantum super Wheatstone bridge with many interesting balance conditions, also containing the standard quantum Wheatstone bridge, where unknown Hamiltonian strengths can be determined by the profiling of heat currents. This circuit can have a plethora of applications in process tomography, metrology, and quantum sensing. A model of a quantum thermal adder circuit for temperature gradients was also constructed, creating further opportunities to design thermal operational amplifiers and other integrated circuits. We also bring out the interrelation between the heat current developed here and the established forms of current used in the literature. Thus, for example, the spin current, which can be experimentally measured [60, 61, 53], provides a handle for experimental determination of the heat current. The framework of the circuits developed here can be adapted to experimental realization. Thus, for example, progress has been made towards the diode architecture with pentamethyl-disilane in an NMR register [62]. Also, an experimental study of the thermodynamic uncertainty relation in the two-qubit NMR spin system, sodium fluorophosphate, was performed in [63], and an experimental setup of a quantum thermal transistor using a superconducting circuit was proposed in [64]. Along similar lines, experimental demonstration of other quantum thermal circuits discussed here could be envisaged. Our analysis opens up several research avenues, such as studies in thermal conductivity, balance conditions of the super Wheatstone bridge, and other analogs of electric circuits. It is thus sufficient to say that the present work offers the fundamentals of a comprehensive theory of quantum thermal analogs of electric circuits, paving the way for the corresponding integrated circuits.
@@ II Details on the forms of heat currents and proof of quantum thermal version of Kirchhoff’s current law
Consider a general system of $n$ nodes (labeled $1$, 22, 33, …, $n$), where each node is coupled to all other nodes. Further, each node is weakly coupled to its respective bosonic thermal bath (labeled $I$, IIII, IIIIII, …, $N$). The system’s Hamiltonian in this setup is given by
@@
where $\mathcal{D}_{jk}(\rho)=\gamma_{j}(\tilde{N}_{jk}+1)\left(L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\left\{L_{k}^{\dagger}L_{k},\rho\right\}\right)+\gamma_{j}\tilde{N}_{jk}\left(L_{k}^{\dagger}\rho L_{k}-\frac{1}{2}\left\{L_{k}L_{k}^{\dagger},\rho\right\}\right)$, and $\tilde{N}_{jk}=\frac{1}{e^{\beta_{j}\omega_{k}}-1}$, with $\beta_{j}=T_{j}^{-1}$ for (j,k)=(I,1),(II,2),,(N,n)(j,k)=(I,1),(II,2),...,(N,n). Further, $L_{k}$ are the Lindblad jump operators, which in the case of qubits are $\sigma^{-}_{k}$. Let us pick a node $m$ (Hamiltonian $H_{m}$) from the $n$ nodes. This node $m$ would have quantum heat currents from all the sources, that is, from all the other nodes that interact with it, as well as from the bosonic bath it is in interaction with (let the corresponding bath be $M$). The time derivative of the expectation value of $H_{m}$ is
@@
It can be easily verified from the above equation that ${\rm Tr}\left[H_{m}\mathcal{D}_{jk}(\rho)\right]=0$ for (j,k)=(I,1),(II,2),,(N,n)(j,k)=(I,1),(II,2),...,(N,n) except when $j=M$ and $k=m$. Therefore, Eq. (18) reduces to
@@
The commutator of $H_{m}$ with $H_{S}$ can further be simplified to give $H_{m},H_{S}]=-\sum_{l=1,l\neq m}^{n}[H_{lm},H_{m}]$, using which we can write
@@ II.1 Comparison of the derived heat current with the established forms in the literature
where $j_{k-1,k}$ is the flux incoming from the k1k-1-th qubit to $k$-th qubit and similarly $j_{k,k+1}$ is the flux outgoing from the $k$-th qubit to $k+1$-th qubit. The left-hand side in the above equation can be obtained using the master equation (15) as
@@
where the terms ${\rm Tr}\left[\sigma^{z}_{k}\mathcal{D}_{A1}\right]$ and ${\rm Tr}\left[\sigma^{z}_{k}\mathcal{D}_{B2}\right]$ become zero for $k\neq 1,2$. Further, the commutator $H_{S},\sigma^{z}_{k}]$ simplifies to $H_{S},\sigma^{z}_{k}]=[H_{k-1,k},\sigma^{z}_{k}]+[H_{k,k+1},\sigma^{z}_{k}]$, where $H_{k-1,k}$ and $H_{k,k+1}$ are the interaction Hamiltonians between qubits at sites k1k-1 and $k$ and between qubits at sites $k$ and $k+1$, respectively. This follows from the fact that all other terms in $H_{S}$ commute with $\sigma^{z}_{k}$. Now, the above equation can be rewritten as
@@
$j_{k,k+1}=2J\left\langle\sigma^{x}_{k}\sigma^{y}_{k+1}-\sigma^{y}_{k}\sigma^{x}_{k+1}\right\rangle=I_{M}.$ (34)
A form of heat current is also found by multiplying the spin current by the frequency of the qubit [58]. However, on comparing the spin current $I_{M}$ and the form of heat current obtained in Eq. (23), we find that the magnitudes of both are related by
@@
The forms of heat current, Eqs. (23) and (24), are more general in the sense that the expressions of heat current, along with their correct frequency scaling factors, can be obtained in scenarios where there are chains of harmonic oscillators or a mixed chain of harmonic oscillators and spins, and for arbitrary interactions between them. Further, there is an advantage in using these forms; for example, in a complicated scenario, such as a circuit of a thermal super Wheatstone bridge, the magnitude and direction of the heat current can be exactly determined.